Surface Area= length=12, width= 17, height=8
Answer by mananth(16946)

there are six surfaces
1 surface = 12*17
2 surface = 12*8
3 surface = 8*17
there are two of each surface
2*12*17 +2*12*8 + 2*8*17
= 872 square units